OpenStudy (iwanttogotostanford):

HELP WITH ACT PRACTICE QUESTION: if f(x)=x^2+x+5 and g(x) = sqrroot x, then what is the value of g(4)/f(1) ?

OpenStudy (freckles):

enter 1 into the expression named f enter 4 into the expression named g

OpenStudy (iwanttogotostanford):

why?? I'm doing a practice test and i got this one wrong so i don't need the answer as much just how to do the problem!:-)

OpenStudy (dinamix):

2/7

OpenStudy (iwanttogotostanford):

@dinamix @freckles can you explain why?

OpenStudy (freckles):

have you entered 1 into the expression named f yet?

OpenStudy (iwanttogotostanford):

@freckles yes

OpenStudy (freckles):

f(x)=x^2+x+5 replace x with 1 g(x)=sqrt(x) replace x with 4

OpenStudy (freckles):

what do you get when you evaluate x^2+x+5 for x=1 and what do you get when you evaluate sqrt(x) for x=4?

OpenStudy (iwanttogotostanford):

7 and 2

OpenStudy (freckles):

correct \[\frac{g(4)}{f(1)}=\frac{\sqrt{4}}{1^2+1+5}=\frac{2}{7}\]

OpenStudy (dinamix):

that correct @iwanttogotostanford

OpenStudy (iwanttogotostanford):

ok thanks that wasn't so hard I guess